You need 50mL of an HCl solution that is 0.02N. You have on hand a 0.5N solution. How would you prepare this solution to give the desired volume and concentration?
1
Expert's answer
2019-11-04T08:07:51-0500

For hydrochloric acid:

C1(HCl) = 0.5N = 0.5M

C2(HCl) = 0.02N = 0.02M


n(HCl) = C2(HCl) × V2(HCl) = 0.02M×0.05L = 0.001mol.

V1(HCl) = n(HCl) /  C1(HCl) = 0.001mol/0.5M = 0.002L = 2mL.


It is necessary to take 2mL of 0.5N hydrochloric acid and bring the volume of the solution with distilled water to 50mL.